Write equation in standard form.
Y+1=x+2

Answer by ashleygurl92(6)

Y+1=x+2
first, you subtract one from both sides to get y by itself.
y=x+1
then you subtract x from both sides because you need x & y on both sides but you have to have x before y because you need it in alphabetical order.
so, your answer is
x+y=1
hope this helps!! :)